Do not round intermediate calculations. Give your final answer(s) to three significant figures. Each of the links A B and CD is made of aluminum (E -10.9 x 10 psi) and has a cross-sectional area of 0.22 in2. Knowing that they support the rigid member B C, determine the deflection of point E. 2.8 kip 16 in. 18 in 9 in. n.
Answer #1

here we have

\sumB=0

2.8*9=Fcd*27

Fcd=0.9333kip

taking sum of forces  in vertical direction

\sumFy=0

Fab*27=2.8*18

Fab=1.8667 kip

so now

\deltaB=1.8667*103*16/10.9*106*.22=12.455*10-3

\deltac=0.9333*103*16/10.9*106*0.22=6.227*10-3

and deflection at E is

\deltaE=\deltac+X

\deltaB-\deltac/X=27/18

(12.455-6.227)*10-3/X=1.5

X=4.152*10-3

\deltaE=\deltac+x=6.227*10-3+4.152*10-3=10.379*10-3in

so the answer is 10.379*10-3in or 0.01 in 0.0104 in or 0.010379 in